OpenStudy (anonymous):

Find the slope of the line 2x+5y=-3.

OpenStudy (anonymous):

solve for y. First subtract 2x from both sides of the equation, 5y=-2x-3, then divided by 5 on both sides of the equation, we'll have y=-2/5 x - 3/5. Slope is the number that next to x, so slope= -2/5

OpenStudy (anonymous):

thanks. can you please help me with one more

OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

same thing, subtract 2x, then we have 5y=-2x-10, then divided by 5 for all the terms. So we have y= -2/5 x - 2. So slope is -2/5